# Code Prompting Elicits Conditional Reasoning Abilities in Text+Code LLMs

### v2

@@ 5.2 Code Syntax Elicits Reasoning Abilities
I. Atomic Statements. Inspired by Min et al. (2023), we transform each NL sentence555We only transform the facts in BGQA since transforming the rules into atomic statements as well yields worse results. into a sequence of atomic statements, which we then append to the original sentence. In this way, the atomic statements can be seen as defining variables for each key entity in the text. Hence, this new prompt would resemble code but without control flow and in natural language form. The prompt retains access to the original instance text (i.e., no loss of information) but is also augmented by simplified sentences in the form of atomic statements. This setup allows us to investigate whether the simplicity of the input triggers improves reasoning abilities, regardless of the text and code syntax.
II. Back-Translated Code. In our second experiment, we investigate whether the semantics of the code statements and not the code syntax are the reason behind the performance boost. For this purpose, we back-transform the code prompts into NL such that the reasoning statements (i.e., the if conditions) are clearly and concisely stated in natural language. Specifically, we map every variable into the format Key entity: variable without snake case. For instance, the variable husband_pass_away from Figure 2 would be back-transformed as Key entity: husband pass away. To transform the if statements, we create a translation prompt by providing four demonstrations. These demonstrations simply translate the conditional statements within the code-formatted instance back into natural language. We also translate the variables in the same manner. This makes the back-translated text as close as possible to the code text. We provide examples of this in Table 11 from Appendix H.
Results. The results666We do not conduct ablation tests on ShARC because, as explained in Section 5, these ablations aim to understand why code prompts outperform text prompts using the highest performing model. in Table 2 show that (1) prompting with atomic statements does not reach the performance of code prompts, and (2) mapping back from code to NL results in a performance drop compared to code prompts. These findings suggest that code prompts enhance LLM performance beyond mere text simplification. This conclusion is supported by the observation that these alternative text simplification approaches, despite offering similar semantics to code prompts, fail to replicate the performance gains observed with code prompts. Therefore, these results imply that specific syntactic features embedded within code directly contribute to performance improvement.
@@ 5.3 Code Semantics are Important
Previously, we have shown that code syntax is necessary to elicit the reasoning abilities of text+code LLMs. Now, we aim to investigate which aspects of code are pivotal. In particular, we evaluate the impact of retaining the natural language text of the original instance within the code comments and the importance of the code semantics. To analyze the former, we have (1) removed the code comments that include the original natural language text from the input and evaluated the performance of the new prompts. To analyze the latter, we (2) perturbed the code to anonymize the variables and functions, as well as (3) added random code whose semantics are completely irrelevant to the original natural language text. In the latter two cases, the code comments remain unmodified (examples illustrating them are provided in Table 12 from Appendix H). Since CondQA includes span answers and removing the NL text would make it impossible for the model to generate the span, we only report performance on the yes-no answers partition (CondQA-YN).
@@
These results confirm that code alone does not trigger reasoning abilities, and instead, the combination of code that represents the original natural language instance and the NL text is able to unlock the potential of LLMs.
| Prompt | CQA | CQA-YN | BG1 | BG2 | BG3 |
| Anonym. | |||||
| Random | |||||
| - Comments | N.A. |
@@ 5.4 Code Prompts are More Sample-Efficient at Eliciting Reasoning Abilities
Figure 3 shows that when we only provide one demonstration per class (i.e., answer type in our datasets), the performance gap is the largest across all datasets. As expected, this gap decreases when we provide more demonstrations. Moreover, we also observe that code prompts with only one demonstration per class even outperform text prompts with three demonstrations per class, which further shows the sample efficiency of code prompts. These results indicate that code prompts trigger conditional reasoning more efficiently than text prompts on GPT 3.5, and this is one of the reasons for its superior performance.
@@ 5.5 Code Prompts Improve Variable Tracking in LLMs
Code Prompting improves State Tracking. From Table 4, we further observe that Text Prompts make significantly more memory errors than code prompts on all datasets. Specifically, the gap is consistently more than 30% with peaks on CondQA (66.69%) and BGQA-3 (39.8%). Therefore, this experiment empirically confirms our hypothesis that code prompts improve state tracking of the key entities and variables when compared to text prompts.
@@ 6 Conclusions and Future Work
This work demonstrates that conditional reasoning abilities can be triggered by using code prompts in large language models (LLMs) of text and code. These code prompts contain the original natural language (NL) formulation as a code comment and code that formulates the logic of the text. To create these code prompts, we use in-context learning to teach an LLM how to conduct such a transformation automatically. Through multiple experiments on three LLMs and three datasets, we show that code prompts trigger conditional reasoning abilities, with large performance gains w.r.t. text prompts (up to percentage points on GPT 3.5, on Mixtral, and Mistral). Our experiments show that even simple code can be beneficial as long as it closely follows the semantics of the NL text and is accompanied by the original NL text. We also show that code prompts are more sample-efficient than text prompts and that their performance boost stems from their superior ability to identify and track the state of key variables or entities, a central aspect of the logical dimension of semantic inference.
In our future work, we plan to extend to a wider range of reasoning abilities, such as multi-hop reasoning, to verify the generalization of our findings. We also plan to conduct experiments investigating how pretraining on text, code, and text+code affects the triggering of reasoning abilities.
@@
Transforming a natural language problem into code requires an intermediate step that raises the cost of the whole pipeline. However, this mapping is not a complicated task, as even the smallest models we considered were able to perform it successfully. Therefore, we believe it would be possible to train a small generative model to do it instead of using a large language model. In this way, we could minimize the cost of using code prompts without affecting its performance.
@@
It is our sincere goal to contribute to the social good in multiple ways. Firstly, we note that large language models are massively being adopted by companies to improve their products. However, it is also known their limitations, such as hallucinations and lack of faithfulness, among others. Our work aims to improve the reasoning abilities of LLMs. Also, the use of code prompts may simplify the explainability of the model responses since we can inspect the entailment status of the variables. We hope these results contribute to enhancing the trustworthiness and safety of LLMs. Nevertheless, every development may pose some risks. In our case, the improvement of the reasoning abilities in LLMs may utilized by malicious actors to propagate more persuasive disinformation.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1), by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE, and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Datasets
@@
is a conversational QA dataset with natural language rules where most questions are underspecified. Therefore, the model may need to ask a follow-up question to know more about the background of the interlocutor to return an answer. The documents are of legal domain retrieved from the web pages of different governments and state agencies. Since this is a conversational QA and we are not interested in evaluating the conversational abilities of LLMs, we transform the task into regular QA, instead of conversational QA. To do this, the model must answer yes, no, or not enough information for each question. In the original task, not enough information, would lead to the generation of a follow-up question. The test set was released less than one week before the deadline for submission of this paper.777https://sharc-data.github.io/index.html Therefore, at the time of experimentation, we randomly divided the dev set into two equal partitions and used one for dev and the other one for test.
@@
| ShARC | 21890 | 1135 | 1135 | CC-BY-SA-3.0 |
Appendix B Prompt Formulation
@@
Firstly, we define the components of a data point in this dataset: facts (), rules (), and questions (). Therefore, our text prompt is defined as follows888BGQA provides a field example with all the variables of the dataset concatenated with descriptions. We use this field as text prompt.:
@@
Appendix C LLM Setup
@@ Appendix D Costs
Appendix E Results on Small LMs with Short Context Window
We have shown the effectiveness of code prompting in the most popular sizes of LLMs in table 1 from section 5.1. However, it is becoming increasingly popular the development of small language models (sLMs) due to their cheaper inferece cost and higher token thoughput (Gunasekar et al., 2023). Therefore, we have conducted a preliminary experiment with Phi-2999https://huggingface.co/microsoft/phi-2, a text+code model of 2.7B parameters on BGQA-1 to show that our prompting methodology also holds in sLMs. As we can show on table 9, code prompting yields a remarkable performance boost of 15 points. However, due to the limited context window of Phi-2, it is not straightforward to conduct in-context learning on our other datasets.
@@ Appendix E Results on Small LMs with Short Context Window
Appendix F Qualitative Analysis
Our previous experiments have focused on the accuracy of the answers. Obtaining a correct answer is correlated to a correct reasoning process. However, it is possible to obtain a correct answer despite an incorrect reasoning chain. Due to the difficulty of automatically evaluating the chain of thoughts, we perform a limited human evaluation. We sample ten instances from BGQA-3 (i.e., the dataset with the most complex reasoning chains) where both text and code prompts return the correct answer and manually inspect the quality of the reasoning chains. Based on this limited sample, we observe that text prompts conduct a 100% correct reasoning chain in only two cases, while code prompts do so in three cases. We identify two main types of errors: (1) wrong conditional reasoning and (2) commonsense errors. We provide examples of those in Table 10. This observation raises the question of whether the generated chain of thought is not faithful to the internal reasoning of the model, as suggested by Lyu et al. (2023) or whether the model generated the right answer from a greedy attempt to reach the closest plausible conclusion (code prompts shows a reasoning error in the last step of the CoT in three out of seven cases). We leave the analysis of faithfulness of chains of thoughts as future work.
| Error Type | Example | Explanation |
Commonsense |
We know the catfish has a harmonica, and according to Rule3 "if the catfish has something to sit on, … |
The model believes a harmonica is something to sit on. |
Conditional Reasoning |
# We know the lion does not remove from the board one of the pieces of the dog, and according to Rule5 "if something becomes an enemy of the squid but does not remove from the board one of the pieces of the dog, then it steals five points from the mosquito." become_enemy(lion, squid)==True not remove_piece(lion, dog)==True steal_points(lion,mosquito,5)=rule5(lion) steal_points(lion, mosquito, 5)==True |
We do not know become_enemy(lion, squid) == True, but if we assume this, we reach the question variable, so we get an answer to the question. |
Appendix G Atomic Statements
Appendix H Examples of Code Ablations
Appendix I Variable Tracking Setup
@@
Code |
# <p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p> if applicant_age >= 18 and entitled_inheritor and closest_relative: can_apply_estate_administrator = True |
Code NL |
<p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p> if you are 18 or over and you are the most entitled inheritor of the deceased’s estate and you are the closest living relative, you can apply to become the estate’s administrator |
Code |
# Rule2: Be careful when something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish because in this case it will surely not burn the warehouse of the mosquito (this may or may not be problematic) rule2(something) = remove(something, piece_of(dog)) & enemy(something, catfish) => not burn(something, warehouse_of(mosquito)) |
Code NL |
Rule2: If something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish, then it does not burn the warehouse of the mosquito |
@@
Original Code |
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if left_country_and_fear_persecution: eligible_for_asylum = True |
Anonymous Code |
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if var_1 var_2 = True |
Random Code |
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if value_of_property_gone_down_by_more_than_50: eligible_to_claim = True getting_housing_benefit = True |
@@
Problem instance |
Human |
Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court? Document: <h1>What is a special guardian</h1> <p>You can apply to be a child’s special guardian when they cannot live with their birth parents and adoption is not right for them.</p> … Answers can be "yes" or "no". Let’s think step by step: |
Partial CoT |
AI |
@@
Probe |
Human |
Now, I want to ask you about the value of some key entities you used. Your answers must be ‘True‘, ‘False‘, ‘unknown‘, or a string. It is very important that you only write the exact value. From the speaker perspective, is it true that the children have been living with me for the last 4 years? |
Probe |
AI |
True |
Appendix J Confusion Matrices
Appendix K Prompt Examples

### v3

@@ 5.2 Code Syntax Elicits Reasoning Abilities
I. Atomic Statements. Inspired by Min et al. (2023), we transform each NL sentence444We only transform the facts in BGQA since transforming the rules into atomic statements as well yields worse results. into a sequence of atomic statements, which we then append to the original sentence. In this way, the atomic statements can be seen as defining variables for each key entity in the text. Hence, this new prompt would resemble code but without control flow and in natural language form. The prompt retains access to the original instance text (i.e., no loss of information) but is also augmented by simplified sentences in the form of atomic statements. This setup allows us to investigate whether the simplicity of the input triggers improves reasoning abilities, regardless of the text and code syntax.
II. Back-Translated Code. In our second experiment, we investigate whether the semantics of the code statements and not the code syntax are the reason behind the performance boost. For this purpose, we back-transform the code prompts into NL such that the reasoning statements (i.e., the if conditions) are clearly and concisely stated in natural language. Specifically, we map every variable into the format Key entity: variable without snake case. For instance, the variable husband_pass_away from Figure 2 would be back-transformed as Key entity: husband pass away. To transform the if statements, we create a translation prompt by providing four demonstrations. These demonstrations simply translate the conditional statements within the code-formatted instance back into natural language. We also translate the variables in the same manner. This makes the back-translated text as close as possible to the code text. We provide examples of this in Table 16 from Appendix P.
Results. The results555We do not conduct ablation tests on ShARC because, as explained in Section 5, these ablations aim to understand why code prompts outperform text prompts using the highest performing model. Results for Mistral and Mixtral are shown in Appendix J. in Table 2 show that (1) prompting with atomic statements does not reach the performance of code prompts, and (2) mapping back from code to NL results in a performance drop compared to code prompts. These findings suggest that code prompts enhance LLM performance beyond mere text simplification. This conclusion is supported by the observation that these alternative text simplification approaches, despite offering similar semantics to code prompts, fail to replicate the performance gains observed with code prompts. Therefore, these results imply that specific syntactic features embedded within code directly contribute to performance improvement.
@@ 5.3 Code Semantics are Important
Previously, we have shown that code syntax is necessary to elicit the reasoning abilities of text+code LLMs. Now, we aim to investigate which aspects of code are pivotal. In particular, we evaluate the impact of retaining the natural language text of the original instance within the code comments and the importance of the code semantics. To analyze the former, we have (1) removed the code comments that include the original natural language text from the input and evaluated the performance of the new prompts. To analyze the latter, we (2) perturbed the code to anonymize the variables and functions, as well as (3) added random code whose semantics are completely irrelevant to the original natural language text. In the latter two cases, the code comments remain unmodified (examples illustrating them are provided in Table 17 from Appendix P). Since CondQA includes span answers and removing the NL text would make it impossible for the model to generate the span, we only report performance on the yes-no answers partition (CondQA-YN).
@@
These results confirm that code alone does not trigger reasoning abilities, and instead, the combination of code that represents the original natural language instance and the NL text is able to unlock the potential of LLMs. We observe similar patterns on Mistral and Mixtral in Appendix J.
Prompt | CQA | CQA-YN | BG1 | BG2 | BG3 | |
Anonym. | ||||||
Random | ||||||
- Comments | N.A. |
@@ 5.4 Code Prompts are More Sample-Efficient at Eliciting Reasoning Abilities
Figure 3 shows that when we only provide one demonstration per class (i.e., answer type in our datasets), the performance gap is the largest across all datasets. As expected, this gap decreases when we provide more demonstrations. Moreover, we also observe that code prompts with only one demonstration per class even outperform text prompts with three demonstrations per class, which further shows the sample efficiency of code prompts. These results indicate that code prompts trigger conditional reasoning more efficiently than text prompts on GPT 3.5, and this is one of the reasons for its superior performance. We conduct additional analysis on Mistral and Mixtral in Appendix K.
@@ 5.5 Code Prompts Improve Variable Tracking in LLMs
Code Prompting improves State Tracking. From Table 4, we further observe that Text Prompts make significantly more memory errors than code prompts on all datasets on GPT 3.5 (Results for Mistral and Mixtral are provided in Appendix M). Specifically, the gap is consistently more than 30% with peaks on CondQA (66.69%) and BGQA-3 (39.8%). Therefore, this experiment empirically confirms our hypothesis that code prompts improve state tracking of the key entities and variables when compared to text prompts.
6 Human Analysis of the Generated Code
We conduct a small human evaluation to confirm the faithfulness of the generated code to the source natural language text. We evaluate the code generations of all our models on ten random samples from the dev set of CondQA, ShARC, and BGQA (in particular, we use BGQA-1 partition). We check for perfect translations, and for the failing cases, we analyze the errors.
BGQA.
We observe perfect translations in all models for all the analyzed samples. We attribute this effectiveness to the close alignment between the natural language documents and first-order logic.
ShARC.
We observe that GPT 3.5 generates perfect translations in all cases except one. However, this case is a corner case where the document is irrelevant to the question, and therefore, there is no answer. Furthermore, the document is only one line. Consequently, the model does not generate code and simply keeps the text as a code comment. In the case of Mixtral 8x7B, we observed perfect code translations for 70% of the samples. One of the failing cases assings as the question variable a variable that is actually from the conversation history, no the quesiton. Another error case exhibits wrong value assingments to some variables. They should be none, but they are assinged true and false. The last case is the corner case explain above. As for Mistral 7B, we find that 60% of the analyzed samples have a perfect translation. In the remaining 40%, we observe three cases with no question variable and the same corner case as before. However, the semantics of the natural language text remain, thanks to the code comments.
CondQA.
We observe that GPT 3.5 generates a perfect translation in 8 out of the 10 cases. In these two cases with errors, we observe that most of the code is correct, but in both cases, one conditional statement is missed. The model directly generates the body of the if statement without the corresponding if. It is also worth noting that the code is of high quality, including data structures as ditionaries, generating code that explains tables, and also generates lists of strings. In the case of Mixtral 8x7B, we obtain perfect translations in 6 out of 10 cases. All the failing cases exhibit the same type of error: there is a variable statement without a prior if condition. It is worth noting that the code, in general, is of high quality, contains data structures such as dictionaries and even process tables. Lastly, in the case of Mistral 7B, we observed a bit worse results. Only 4 out of the 10 cases are perfect translations. In two cases, there is no code, and instead, the model only generated the original natural language text in code comments. We also observe one case where the first half of the text is correctly translated into code but the second half only contains the code comments representing the natural language text. We also observe one case where the code is correct, but the indentation is wrong; all code blocks are under the first if statement, which should not be like that. Lastly, we find two cases where the if statements do not contain an execution body. Notably, even in the cases where the code is not perfect, the original semantics from the natural language remain untouched because they are preserved through code comments.
This analysis confirms that, in general, the translated code faithfully represents the semantics of the source natural language text.
@@ 7 Conclusions and Future Work
This work demonstrates that the code representation of a natural language task (i.e., code prompts) can elicit reasoning abilities in large language models (LLMs) of text and code. These code prompts contain the original natural language (NL) formulation as a code comment and code that formulates the logic of the text. To create these code prompts, we use in-context learning to teach an LLM how to conduct such a transformation automatically. Through multiple experiments on three LLMs and three datasets, we show that code prompts trigger conditional reasoning abilities, with large performance gains w.r.t. text prompts (up to percentage points on GPT 3.5, on Mixtral, and Mistral). Our experiments reveal that even simple code can be beneficial as long as it closely follows the semantics of the NL text and is accompanied by the original NL text. We also show that code prompts are more sample-efficient than text prompts and that their performance boost stems from their superior ability to identify and track the state of key variables or entities, a central aspect of the logical dimension of semantic inference.
In our future work, we plan to extend to a wider range of reasoning abilities, such as multi-hop reasoning, to understand the capacity and generalizability of code prompting. We also plan to investigate how pretraining on text, code, and text+code affects the triggering of LLMs’ reasoning abilities.
@@
Transforming a natural language problem into code requires an intermediate step that raises the cost of the whole pipeline. However, this mapping is not a complicated task, as even the smallest models we considered were able to perform it successfully in an in-context learning setup. Therefore, we believe it would be possible to train a small generative model to do it instead of using a large language model. In this way, we could minimize the cost of using code prompts without affecting its performance.
@@
This work focuses only on analyzing the effects of code representations for natural language tasks. However, it could be possible that other input representation spaces also elicit reasoning abilities. We limit the scope of this work to only the space of simple structured languages (more details on Appendix D) because prior research suggests that pretraining on code improves the reasoning abilities of LLMs, but it might be possible that certain natural languages such as German or Chinese, or certain types of programming languages, such as declarative or logical, also elicit certain abilities. Similarly, we do not conduct experiments on multiple code generation methods because our goal is to analyze whether the mere change of the representation can elicit reasoning abilities and not an analysis of the best coding style.
Our tasks require instruction following abilities, so we do not conduct comparisons of base vs. chat models. Future work could investigate whether instruction tuning has an impact on the LLMs’ abilities to understand code.
@@
Our work aims to improve the reasoning abilities of LLMs. The use of code prompts may also simplify the explainability of the model responses since we can inspect the entailment status of the variables. We hope these results contribute to enhancing the trustworthiness and safety of LLMs. Nevertheless, every development may pose some risks. In our case, the improvement of the reasoning abilities in LLMs may be utilized by malicious actors to propagate more persuasive disinformation.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1), by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE, and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Somak Aditya acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under grant agreement No 951847.
We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
Lastly, we thank Max Glockner, Jonathan Tonglet, Sheng Lu, and the anonymous reviewers for their insightful comments and suggestions on a prior draft of this paper.
Appendix A List of Appendices
We start the appendices with details on the datasets and prompts (Appendix B and C). Then, we provide additional details on the features of the generated code, the setup of the LLMs, and their costs (Appendix D, E, and F). Subsequently, we perform experiments on code-only and text-only LLMs to show that code prompting should only be used on text+code LLMs (Appendix G and H). Then, we show experiments on small LLMs (Appendix I), ablations on local LLMs (Appendix J, K, L, and M. We conclude with the confusion matrices for the main experiments Appendix N, prompt examples O, P, and Q.
Appendix B Datasets
@@
is a conversational QA dataset with natural language rules where most questions are underspecified. Therefore, the model may need to ask a follow-up question to know more about the background of the interlocutor to return an answer. The documents are of legal domain retrieved from the web pages of different governments and state agencies. Since this is a conversational QA and we are not interested in evaluating the conversational abilities of LLMs, we transform the task into regular QA, instead of conversational QA. To do this, the model must answer yes, no, or not enough information for each question. In the original task, not enough information, would lead to the generation of a follow-up question.
@@
ShARC | 21890 | 2270 | 8276 | CC-BY-SA-3.0 |
Appendix C Prompt Formulation
@@
Firstly, we define the components of a data point in this dataset: facts (), rules (), and questions (). Therefore, our text prompt is defined as follows666BGQA provides a field example with all the variables of the dataset concatenated with descriptions. We use this field as text prompt.:
@@
Appendix D Coding Features
Appendix E LLM Setup
@@
All of our prompting methods are implemented using the Langchain library.777https://github.com/langchain-ai/langchain We set the decoding temperature to zero and use greedy sampling to make the outputs deterministic. For each experiment, we use a random sample from the training set as demonstrations. The LLM generating the code for code prompts is the same one as the one running the code to generate the final answer. We evaluate each model and prompt in the dev set of each dataset with two random seeds. Since the demonstrations are selected randomly, the seed determines them. The seed that yields the best performance on the dev set is then used for the final evaluation on the test set.
@@
We use chain of thoughts (CoT) based on the provided annotations of the datasets. We do not use advanced CoT methods for text prompts because our aim is to quantify how much improvement we can get by transforming the natural language CoT into code syntax, and therefore, the natural language text and code must be as close as possible. The use of advanced CoT methods would also be reflected in the code syntax, making the experimental setup more complicated without providing better insights.
@@ Appendix F Costs
Appendix G Code-only LLMs
[…]
However, we observe a different behavior on BGQA. In this dataset, code prompts outperform text prompts. We attribute this to the high alignment with the first-order logic of this dataset, which makes it closer to the intended use of the model.
Nevertheless, it is important to note that these results are not intended to be comprehensive enough to conclude that code LLMs or Code Llama can or cannot solve natural language tasks, which is out of the scope of this work. Instead, they simply seem to confirm the warnings of the authors of Code Llama, i.e., this model is not intended for natural language tasks.
Model | Text Prompt | Code Prompt |
CondQA | 31.58 | 26.34 |
ShARC | 58.33 | 18.62 |
BGQA1 | 44.38 | 44.78 |
BGQA2 | 44.59 | 49.41 |
BGQA3 | 40.88 | 46.44 |
Appendix H Text-only LLMs
As briefly mentioned in Section 4.3, text-only LLMs are not expected to perform well with code prompting and should not be used for this as they are not explicitly trained on code. For example, on the MBPP coding benchmark (Austin et al., 2021), LLaMA 2 scores 26.1% , while Mistral 7B (a text+code LLM) achieves 47.5% and code-llama 2 7B, achieves 52.5% (Jiang et al., 2023). Table 10 further proves our claim. LLaMA 2 7B-Chat (Touvron et al., 2023b) with code prompting never outperforms text prompting due to its lack of code understanding.
Prompt | CQA | ShARC | BG1 | BG2 |
Text | 29.79 | 46.16 | 51.85 | 39.23 |
Code | 21.32 | 24.74 | 45.16 | 37.66 |
Appendix I Results on Small LMs with Short Context Window
We have shown the effectiveness of code prompting in the most popular sizes of LLMs in table 1 from section 5.1. However, it is becoming increasingly popular the development of small language models (sLMs) due to their cheaper inferece cost and higher token thoughput (Gunasekar et al., 2023). Therefore, we have conducted a preliminary experiment with Phi-2888https://huggingface.co/microsoft/phi-2, a text+code model of 2.7B parameters on BGQA-1 to show that our prompting methodology also holds in sLMs. As we can show on table 11, code prompting yields a remarkable performance boost of 15 points. However, due to the limited context window of Phi-2, it is not straightforward to conduct in-context learning on our other datasets.
@@ Appendix I Results on Small LMs with Short Context Window
Appendix J Ablations on Local LLMs
Table 12 shows that all ablations to code prompting in Mistral yield significant performance drops similar to those observed in GPT 3.5
Prompt | BG1 | BG2 | BG3 |
Atomic St. | -4.29 | -8.09 | -2.19 |
NLCode | -2.7 | -4.36 | -5.97 |
Anonym. Code | -0.45 | -7.17 | -4.52 |
Rnd Code | -4.27 | -11.16 | -13.93 |
- Comments | -0.79 | -9.88 | -29.21 |
Table 13 shows that most ablations to code prompting in Mixtral also yield significant performance drops similar to those observed in GPT 3.5, except on BGQA-1, where text prompts outperform code prompts. Therefore, it is expected that the atomic statements and natural language code ablations improve performance.
Prompt | BG1 | BG2 | BG3 |
Atomic St. | 1.47 | -4.01 | -7.61 |
NLCode | 0.9 | -2.82 | -3.18 |
Anonym. Code | 0.21 | -16.59 | 1.25 |
Rnd Code | -15.88 | -9.56 | -11.31 |
- Comments | -12.81 | -7.14 | 5.38 |
Appendix K Number of Demonstrations on Local LLMs
We conduct experiments with Mistral and Mixtral on BGQA because it is the dataset where we clearly see a difference between text and code prompts with different numbers of demonstrations. Table 14 shows that Mistral achieves the best results with just one demonstration for both prompting methods. We attribute this to the long length and complexity of the demonstrations999We cannot even provide three demonstrations for BGQA-2 and 3 due to the length of each demonstration., which can confuse LLMs, especially those that are small, which is the case for Mistral. Li et al. (2024) shows that Mistral dramatically decreases its performance as the number of input tokens increases. However, very large LMs are more resilient to the input length. Text prompts on Mixtral achieve the best results with at least two demonstrations, while code prompts only achieve significantly better results with more than one demonstration in BGQA-3.
Model | Prompt | # Dem. | BG1 | BG2 | BG3 |
Mistral | Text | 1 | 48.26 | 45.90 | 48.63 |
2 | 45.86 | 48.82 | 47.92 | ||
3 | 44.25 | N.A. | N.A. | ||
Code | 1 | 51.54 | 55.28 | 52.83 | |
2 | 51.85 | 50.51 | 48.66 | ||
3 | 39.58 | N.A. | N.A. | ||
Mixtral | Text | 1 | 57.06 | 36.71 | 30.8 |
2 | 65.49 | 39.83 | 33.23 | ||
3 | 63.73 | N.A. | N.A. | ||
Code | 1 | 61.85 | 44.93 | 35.22 | |
2 | 58.44 | 45.70 | 39.43 | ||
3 | 61.3 | N.A. | N.A. |
Appendix L Variable Tracking Setup
@@
Appendix M Memory Errors on Local LLMs
Analyzing the memory errors in Mixtral and Mistral becomes much more challenging than in GPT 3.5 due to the lower performance of the models. Table 15 shows that, in general, Mixtral makes significantly more memory errors than GPT 3.5. On BGQA-1, we observed more memory errors on the questions where the model fails than when the model returns the correct answer, as expected. We further see that code prompts make fewer memory errors than text prompts, confirming the results of GPT 3.5. However, the margins are much narrower than in GPT 3.5.
On BGQA-2 and 3, the model surprisingly makes fewer memory errors on the wrong answer partition. However, in both cases, the percentage is so high that makes the test unreliable, as we observe in CondQA in GPT 3.5. In the specific case of Mistral, we were not able to run this experiment successfully due to the model not understanding the task. We believe our experiment setup for memory error analysis only works well on very high-performing models such as GPT. To analyze them on smaller models such as Mistral and Mixtral, we would need to devise other types of experiments, which is out of the scope of this work. We encourage further research on entity tracking methods to analyze LLMs and leave this as future work.
Correct Ans. | Incorrect Ans. | |||
Text | Code | Text | Code | |
BGQA-1 | 58.5 | 54.0 | 60.4 | 55.6 |
BGQA-2 | 70.8 | 77.6 | 66.0 | 80.0 |
BGQA-3 | 65.2 | 80.4 | 62.8 | 94.4 |
Appendix N Confusion Matrices
Appendix O Atomic Statements
Appendix P Examples of Code Ablations
@@
Code | # <p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p> if applicant_age >= 18 and entitled_inheritor and closest_relative: can_apply_estate_administrator = True |
Code NL | <p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p> if you are 18 or over and you are the most entitled inheritor of the deceased’s estate and you are the closest living relative, you can apply to become the estate’s administrator |
Code | # Rule2: Be careful when something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish because in this case it will surely not burn the warehouse of the mosquito (this may or may not be problematic) rule2(something) = remove(something, piece_of(dog)) & enemy(something, catfish) => not burn(something, warehouse_of(mosquito)) |
Code NL | Rule2: If something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish, then it does not burn the warehouse of the mosquito |
@@
Original Code | # <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if left_country_and_fear_persecution: eligible_for_asylum = True |
Anonymous Code | # <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if var_1 var_2 = True |
Random Code | # <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p> if value_of_property_gone_down_by_more_than_50: eligible_to_claim = True getting_housing_benefit = True |
@@
Problem instance | Human | Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court? Document: <h1>What is a special guardian</h1> <p>You can apply to be a child’s special guardian when they cannot live with their birth parents and adoption is not right for them.</p> … Answers can be "yes" or "no". Let’s think step by step: |
@@
Probe | Human | Now, I want to ask you about the value of some key entities you used. Your answers must be ‘True‘, ‘False‘, ‘unknown‘, or a string. It is very important that you only write the exact value. From the speaker perspective, is it true that the children have been living with me for the last 4 years? |
Probe | AI | True |
Appendix Q Prompt Examples